OpenStudy (anonymous):

The radius of convergence for the power series (x-3)^2n/n from n=1 to infinity is equal to 1. What is the interval of convergence?

OpenStudy (zarkon):

what do you think it is?

OpenStudy (anonymous):

well i thought it was from -1 to 1 but its wrong. I know the anwer but just not how to get it

OpenStudy (zarkon):

it has to be centered at 3

OpenStudy (anonymous):

I was trying to use the ratio test but i guess i did something wrong

OpenStudy (zarkon):

not zero

OpenStudy (anonymous):

why 3?

OpenStudy (zarkon):

with the ratio test you will end up with |x-3|<1

OpenStudy (anonymous):

can pleas show me how 2 do the ratio test. i can usually do the probem from there. it ust teh test i have trouble with

OpenStudy (zarkon):

hold on....is this your series... \[\sum_{n=1}^{\infty}\frac{(x-3)^{2n}}{n}\]

OpenStudy (anonymous):

yes andthats equal to 1

OpenStudy (zarkon):

ok...so you still get |x-3|<1 \[\Rightarrow 2<x<4\]

OpenStudy (zarkon):

the radius of convergence is 1...not the sum

OpenStudy (anonymous):

oh right...sorry. Can u show me how u got x-3 less than 1

OpenStudy (zarkon):

ratio test...

OpenStudy (zarkon):

let \[a_n=\frac{(x-3)^{2n}}{n}\] then \[\left|\frac{a_{n+1}}{a_n}\right|\] \[=\left|\frac{\frac{(x-3)^{2(n+1)}}{n+1}}{\frac{(x-3)^{2n}}{n}}\right|\] simplify this and take the limit as \(n\to \infty\)

OpenStudy (anonymous):

i understand how to get that but i dont ever get the simplying part right. im not sure how to cancel out things.

OpenStudy (zarkon):

ratio test says that you will have convergence if the above limit is less than 1

OpenStudy (zarkon):

\[=\left|\frac{\frac{(x-3)^{2(n+1)}}{n+1}}{\frac{(x-3)^{2n}}{n}}\right|\] \[=\left|\frac{(x-3)^{2(n+1)}}{n+1} \frac{n}{(x-3)^{2n}}\right|\] \[=\left|\frac{(x-3)^{2n+2}}{n+1} \frac{n}{(x-3)^{2n}}\right|\] \[=\left|\frac{(x-3)^{2}}{n+1} n\right|\] \[=\left|(x-3)^{2}\frac{n}{n+1} \right|\to|x-3|^2\text{ as }n\to\infty\]

OpenStudy (zarkon):

then \[|x-3|^2<1\Rightarrow |x-3|<1\]

OpenStudy (anonymous):

right so then u add 3 and u get x betwen 2 and 4. so now we have to test them to see which is included

OpenStudy (zarkon):

yes...you now need to test the boundary

OpenStudy (zarkon):

you sould be careful when finding the interval \[|x-3|<1\] \[X-3<1\] and \[-(x-3)<1\] first one \[X-3<1\Rightarrow x<1+3=4\] for the second \[-(x-3)<1\Rightarrow x-3>-1\Rightarrow x>3-1=2\] thus \[2<x<4\]

OpenStudy (anonymous):

so b/c its equal to a number greater than 1 it diverges

OpenStudy (anonymous):

sorry but i dont understand the difference btwn X and x

OpenStudy (zarkon):

plug 2 and 4 gack into the original series for x and see if they converge

OpenStudy (zarkon):

just a typo...should all be x

OpenStudy (anonymous):

oh ok i see....THANKS SOO MUCH!

OpenStudy (zarkon):

good